OpenStudy (anonymous):

please help I only have a few mins Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______ Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3. Numerical Answers Expected!

OpenStudy (anonymous):

If\[f(x)=\frac{ 1 }{ 3 }x-2 ,g(x)=\frac{ 1 }{ 3 }x+3\]replace those into your question and solve for k\[f(x)=g(x)+k \rightarrow \frac{ 1 }{ 3 }x+3=\frac{ 1 }{ 3 }x-2+k\]

OpenStudy (anonymous):

You should then proceed to simplify and solve for k, subtracting the x/3 from both sides\[3=k-2\]i trust you can finish it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks I got it. thanks sssooo much

OpenStudy (anonymous):

:)